OpenStudy (anonymous):

Part A: The sun produces 3.9 ⋅ 1033 ergs of radiant energy per second. How many ergs of radiant energy does the sun produce in 3.25 ⋅ 103 seconds? (5 points) Part B: Which is the more reasonable measurement of the distance between the tracks on a railroad: 1.435 ⋅ 10−3 mm or 1.435 ⋅ 103 mm? Justify your answer

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

Should that say The sun produces \( 3.9 ⋅ 10^{33} \) ergs of radiant energy per second

OpenStudy (anonymous):

yes @jayzdd

OpenStudy (anonymous):

$$\large \sf \frac{3.9 ⋅ 10^{33} ~ergs} { second} \times 3.25 ⋅ 10^3 ~ seconds$$

OpenStudy (anonymous):

ok is there anything else to that?

OpenStudy (anonymous):

@jayzdd ^

OpenStudy (anonymous):

now we get out our calculators :)

OpenStudy (anonymous):

alright so i calculate that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

alright hold on

OpenStudy (anonymous):

i dont know if this is right but i got 1.2675 × 10^30

OpenStudy (anonymous):

close

OpenStudy (anonymous):

$$ 1.2675 × 10^{37}$$

OpenStudy (anonymous):

ohhh ok is that the final answer?

OpenStudy (anonymous):

yes and you can round it

OpenStudy (anonymous):

ok thank you, can you help with part B also please?

OpenStudy (anonymous):

which measurement seems more likely?

OpenStudy (anonymous):

um 1.435 ⋅ 10^3 mm?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

1 mm is very small , so 10^-3 mm is extremely tiny

OpenStudy (anonymous):

right thank you! i have one more question that i have no idea how to work on, can you help me?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Find the value of the following expression: (2^8 ⋅ 3^−5 ⋅ 6^0)^−2 ⋅ 3 to the power of negative 2 over 2 to the power of 3, whole to the power of 4 ⋅ 2^28 (5 points) Write your answer in simplified form. Show all of your steps

OpenStudy (anonymous):

????

OpenStudy (anonymous):

Let's rewrite this long expression

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how do i re write it

OpenStudy (anonymous):

alright so can you help me solve it

OpenStudy (anonymous):

\[ (2^8 ⋅ 3^{-5} ⋅ 6^0)^{−2} ⋅ \frac{3^{-2}}{2^3} 4 ⋅ 2^28\]

OpenStudy (anonymous):

alright so the one is parentheses are first correct?

OpenStudy (anonymous):

I am not sure about the expression 'whole'

OpenStudy (anonymous):

the 4 by the fraction is in square root

OpenStudy (anonymous):

got it?

OpenStudy (spring98):

Is there any answer to the last one?